# Evaluating Generative Ad Hoc Information Retrieval

### v2

@@ 3.1. Preliminary Considerations
In traditional retrieval, the user is presented with a ranked list of documents (list SERP), each typically referenced by a linked title, snippet, and URL. In generative IR, instead, a response text is presented (text SERP), i.e., a sequence of statements, each optionally referencing one or more sources of evidence in support of the statement. A statement can be any consecutive passage of text, ranging from a single word or phrase to a sentence and even one or more paragraphs. In this context, statements are considered ‘atomic’ in the sense that we disregard the nesting of statements of different lengths, and that they support one or more claims that are pertinent to the user’s information need. They are comparable to the concept of ‘atomic/semantic content units’ (Liu et al., 2023a; Nenkova et al., 2007) in summarization evaluation, or ‘information nuggets’/‘retrieval unit’ in traditional IR (Dang and Lin, 2007; Sakai et al., 2011; Sakai, 2023; Chen et al., 2023). A statement can be referencing none, one, or more than one source. References explicitly link to a source, like a web document containing the information on which the generated statement is based and by which it is grounded. The evaluation commences ad hoc, i.e., with a single-query and without session-based or conversational elements.
@@
To estimate the effectiveness of retrieval systems, offline evaluation within a Cranfield-style evaluation setting (Cleverdon, 1967) is the de facto standard in IR research. It is used to estimate the satisfaction of users with the output of a system for a given topic (query) by relying on a pool of documents judged by assessors for a given topic set (Sanderson, 2010). This pool of relevance judgments can be reused in subsequent experiments by matching retrieved document identifiers with the ones in the pool. Once a pool of relevance judgments has been compile for a given topic, large-scale evaluations of alternative search systems can be performed. However, since the output documents of a generative retrieval system are novel at query time, they cannot be straightforwardly matched with existing pools of relevance judgments. This bears similarities to the unjudged document problem, which is traditionally solved by assuming non-relevance (Fröbe et al., 2023)—a heuristic, which is not feasible for generative IR. Therefore, more sophisticated transfer methods are required to adapt offline evaluation for generative retrieval.
@@
Alternatively, evaluation of generative retrieval systems can be conducted in an online fashion (Sallam et al., 2023). Here, for each run of a system configuration, all its output documents are judged anew, without relying on previous judgments, by collecting user feedback about a system (Kelly, 2009), e.g., by rating their satisfaction. This maximizes the manual effort required to judge runs during experimentation, often happens uncontrolled, is expensive and time-consuming to conduct, and is challenging to replicate, repeat, and reproduce (Renaud and Azzopardi, 2012). To mitigate these problems especially in an academic setting, where access to human user data is often limited, much research went into simulating agents to analyze (interactive) information systems (Maxwell et al., 2015; Maxwell and Azzopardi, 2016; Câmara et al., 2022). However, these agents cannot yet compete with “real” human feedback. Automatic evaluations, where the output of one model is judged by another, has been proposed as a possible way forward (Liu et al., 2023b; Yue et al., 2023), but judging the output of generative models by means of other models has itself been criticized (Sakai, 2023; Bauer et al., 2023; Faggioli et al., 2023).
3.2. Components of the User Search Process
@@
To derive suitable evaluation objectives, first, we have to consider the search process a user undergoes when performing an ad hoc search. Specifically, the synthetic search task enabled by generative retrieval systems should be reflected here. Based on Vakkari (2016), a users’ process encompasses four steps: search formulation, source selection, source interaction, and synthesis of information. Each of these can be mapped to capabilities of generative retrieval systems.
@@
First, during Formulation, the user crafts a specific query that expresses the desired search outcome, addressing their information need. This is no different in generative retrieval systems, though what is called a ‘query’ in IR is a ‘prompt’ in artificial intelligence research. To avoid confusion, we stick to the term ‘query’. In this paper, we leave this step entirely to the user who (iteratively) adapts their search formulation. Yet, we do acknowledge that this task may also be framed as a system task with the goal of enhancing the users’ original query with more context or prompt templates, akin to query suggestion and query expansion in traditional retrieval.
Second, during Selection, the user is presented with a result list and can then examine each entry, possibly through surrogates like snippets. They can assess whether the presented results and their information need align and thus build a focused selection of sources. In generative retrieval systems, this stage corresponds to the system selecting sources that contain potentially relevant information.
Third, during Interaction, the user analyzes the content of each previously selected result in-depth. The aim is to extract and structure the relevant information from each source that addresses the knowledge gap that induces their information need. In generative retrieval systems, this step is supported by the model attending to relevant pieces of information previously retrieved.
Finally, during Synthesis, the user assembles the search outcome. They combine relevant information identified in multiple sources into a coherent answer to their query. In generative IR, this corresponds to the inference of the response text addressing all aspects of the user’s query with information from the previously selected sources. This is key in enabling the synthetic search task. Note that interaction and synthesis often commence concurrently.
@@ 3.3. Evaluation Objectives
For each of the components of the search process, we define a corresponding evaluation objective in the context of generative IR. These are not considered evaluation steps, but rather objectives of the evaluation of the system as a whole.
@@
Formulation is reflected in the evaluation of the models’ input prompt. While search formulation is an important component to evaluate, we believe it is out of the scope of this paper, since, as previously argued, the formulation step is left to the user. For further reading, the issue of prompt engineering as an emergent field of research is covered in extensive related work (Shin et al., 2020; Reynolds and McDonell, 2021; Gao et al., 2021; Liu and Chilton, 2022; Sorensen et al., 2022; White et al., 2023; Yang et al., 2023).
@@
Selection is reflected in the assessment of the context from which a generative retrieval system draws its information. The retrieved sources (and any relevant information that was not retrieved) directly impact the quality of the generated response. Therefore, the retrieval objective assesses a system’s ability to identify source documents satisfying a user’s information need. This includes its ability to select (1) relevant (aligning with the users’ information need), (2) diverse (covering a variety of information), (3) informative (containing valuable information), and (4) correct (providing accurate information) documents from a collection.
@@
Mimicking Interaction, generative ad hoc IR models draw upon reference documents as evidence to generate a response. Yet, grounded text generation may suffer from hallucinations of broadly two types (Maynez et al., 2020): intrinsic hallucinations, where the model wrongly modifies information from the source documents, and extrinsic hallucinations, where the model generates information that is not present in the source documents. Both negatively impact the quality of the generated response (Maynez et al., 2020; Lux et al., 2020). Therefore, the grounding objective assesses a system’s ability to correlate its generated output with information from source documents. This includes its ability to (1) identify (find relevant information), (2) paraphrase (restate that information correctly), and (3) establish consistency (not produce contradictions to other sources).
@@
The relevant information across multiple documents has to be synthesized into a single search outcome. This resembles multi-document summarization. Therefore, the presentation objective assesses a systems’ ability to convey information to a user through the generated response in a useful manner, i.e., its ability to produce text that is (1) concise (at a level of granularity sensible given the topic or needed by the user (Dang, 2005)), (2) coherent (in a uniform style), and (3) accessible (written in an understandable way, which, again, is dependent on user needs).
@@ 3.4. Components of the User Model
Generative IR poses a challenge for developing a user model. As it is a new IR paradigm, little to no user feedback, A/B tests, laboratory studies, or user behaviour data is available in the academic context from which insights about user behavior may be derived. Further, the information search process that user behavior is traditionally grounded in is replaced (in part) by the generative system. Additionally, the assumptions of traditional user models are made for list SERPs and thus have to be revisited, taking into account the previously established evaluation objectives. To this end, we contribute a user model for generative IR, extrapolating from established evaluation practices in related fields, like question answering, summarization, and traditional IR. We follow the considerations of Carterette (2011), who argues that an IR-focused user model is constituted by three distinct (sub-)models: (1) a utility model (how each result provides utility to the user), which induces a gain function; (2) a browsing model (how the user interacts with results), which induces a discount function; and (3) an accumulation model (how the individual utility of documents is aggregated), combining the individual gain and discount values.
3.4.1. Utility Model for Generative IR
We first motivate a utility model by surveying literature on evaluation in IR and related fields. We identify 10 dimensions of utility applicable to the ad hoc synthetic search task. These are grouped into five top-level categories of Coherence, Coverage, Consistency, Correctness and Clarity. We further distinguish the unit from which gain is derived, as either an individual statement in the response (statement-level) or the response as a whole (response-level). Figure 5 summarizes the proposed dimensions of utility as taxonomy, divided into response- and statement-level dimensions with corresponding objectives marked.
@@ 3.4.1. Utility Model for Generative IR
Coherence is a response-level dimension of utility and refers to the manner in which the response is structured and presented. This includes arranging statements to form a coherent narrative without contradictions (Radev and McKeown, 1998; Shah et al., 2021) (Logical Coherence), but also a uniform style of speech (Stylistic Coherence), rendering it readable and engaging (Jin et al., 2020; Capra and Arguello, 2023). Both implement the presentation objective at response level, amounting to “Is the response structured well?” (Logical Coherence) and “Does the response have a uniform style of speech?” (Stylistic Coherence).
@@
Coverage measures the cumulative extent to which presented information is pertinent to the users’ information need. It can be subdivided into two forms (Cambazoglu et al., 2021): Broad Coverage, i.e., whether the response covers a breadth of diverse information (Zheng et al., 2012), and Deep Coverage, i.e., whether the response provides in-depth detailed information with high informativeness (Maxwell et al., 2017). Coverage implements the retrieval objective at response level, amounting to “Does the response cover diverse information?”(Broad Coverage) and “Does the response offer detailed information?” (Deep Coverage).
@@
A commonly observed problem with source-based text generation is inconsistency (Huang et al., 2021) between source and generated text, which is detrimental to utility. Inconsistencies may also occur across multiple statements within a response, rendering it both a statement-level and response-level dimension. We refer to the first as Internal Consistency (response level), which involves assessing the consistency between statements that constitute the response, ensuring that they do not contradict each other (Nishino et al., 2019; Sakai, 2023; Capra and Arguello, 2023). It should be noted that this does not mean that different conflicting perspectives on a topic can not be reflected in the response, however, these should be explained. The second, External Consistency (statement level), involves assessing the consistency between a statement and its source document(s), ensuring that the generated text aligns in terms of content and context (Maynez et al., 2020; Yue et al., 2023; Sakai, 2023). External inconsistencies are often introduced through model hallucinations (Ji et al., 2023). Consistency is different from factual correctness, as it only assesses the alignment of a statement with the source, and not its objective truth. Both notions implement the grounding objective but on different levels, amounting to “Is the response free of contradictions?” (Internal Consistency) and “Is the statement conveying from sources accurately?” (External Consistency)
@@
Correctness gauges to which degree the information provided in the response is factual, reliable, and addressing the user’s information needs. We subdivide correctness into Factual and Topical Correctness. The former captures the degree to which a statement reproduces information that can be assumed as objectively true. Yet, outside of small-scale domain-specific evaluation studies (Sallam et al., 2023) fact-checking remains a challenge (Nakov et al., 2021) and is thus often reduced to a simpler approach, framing it in terms of verifiability (Liu et al., 2023b), not truth. Here, the main requirement is that a piece of information can be attributed to a reliable reference, bestowing it correctness (Wikimedia Foundation, 2023; Yue et al., 2023). Topical correctness denotes whether a statement aligns with the users’ information need (Maddalena et al., 2017; Yang, 2017; Roitero et al., 2018). Both operationalize the retrieval objective at the statement level, amounting to “Does the statement state things that are verifiably true?” (Factual Correctness) and “Does the statement state things within the scope of the user’s information need?” (Topical Correctness).
@@
The response given by a generative retrieval system should be expressed in a clear and understandable manner (Zhu et al., 2009; Sameki et al., 2016). This includes the use of language in a concise (Dang, 2005; Sakai, 2023), comprehensible (Cambazoglu et al., 2021), lexically and grammatically correct, and accessible way to the user (Language Clarity). Note that language clarity does not reflect fluency, which is assumed already at human-level for model-generated text (Sakai, 2023), but rather the response being in the appropriate language register. For example, a technical query might warrant an academic style of writing in the response, while a joke question might afford a more jovial tone. Orthogonal to this, the way a statement is written should always clearly communicate the most salient information (Schuff et al., 2022), and where it stems from (Nourani et al., 2019), in order to make the response explainable (Content Clarity). Both operationalize the presentation objective at the statement level, amounting to “Is a statement written in an easily readable way?” (Language Clarity) and “Does the statement put its focus on the most salient points?” (Content Clarity).
3.4.2. Reading Model for Generative IR
@@
For list SERPs, user interaction is modeled by a browsing model, of which two fundamental kinds exist. The set-based kind assumes that a user indiscriminately examines all retrieved documents, while the ranking-based model assumes a user traversing retrieved documents ascending in rank, stopping when either their information need is fulfilled or the search is aborted (Carterette, 2011). Aborting the search is primarily motivated by the effort being too high to justify continuing to browse. Yet, in generative IR, the selection and interaction steps of the search process are supported by the system, thus the user only has to read the generated text. This reduces the effect of stopping criteria grounded in effort, with most users only aborting their search once their knowledge gap is fulfilled, the response is deemed insufficient, or the whole response was read. This is neither set-based, as reading the response is a sequential process and early stopping might occur, nor traditionally ranking-based, as aborting the search is not motivated by effort but rather by search (dis-)satisfaction.
@@
We therefore propose a reading model for generative IR, as an evolution of the standard browsing model, which instead models the attention a user places on each statement while reading. Since there are no empirical studies on reading behavior for generative retrieval at present, we instead turn to related work in reading behavior for document comprehension. We identify a total of six criteria which influence the reading process of documents for an information-seeking purpose, three of which we deem relevant to the case of generative IR. First, Progression (Buscher et al., 2012; Li et al., 2018, 2019; Zheng et al., 2019; Wu et al., 2023) implies that users parse a document sequentially, i.e., progress through the statements constituting the text in order. Second, Decay (Frey et al., 2013; Li et al., 2018, 2019; Zheng et al., 2019; Wu et al., 2023) implies that the reading attention diminishes over the span of the text. Third, Saturation (Li et al., 2018, 2019) implies that users abort once they read enough to fulfill their information need. In sum, this characterizes the browsing behavior of a user as sequentially reading with decaying attention, stopping early if saturated.
@@
While three other characteristics of reading behavior have additionally been found in related work, we deem them superfluous for this reading model: (1) Although perceived relevance is heightened following a relevant statement (Li et al., 2018; Zheng et al., 2019), we adopt the same restriction as a static browsing model for ad hoc retrieval evaluation that neglects inter-statement effects (Moffat et al., 2013, 2015). While the effect is acknowledged, its effect size may not justify the costs of its operationalization. (2) Although attention is highest around query terms (Li et al., 2018; Zheng et al., 2019), we model utility not per token, but at the statement level, thus rendering this effect constant. (3) Although users skip content non-relevant to them (Li et al., 2018; Gwizdka, 2014; Buscher et al., 2012), such statements already receive zero utility.
The properties of the proposed reading model can be related to the (Moffat et al., 2017) framework of browsing models for list SERPs. The conditional ‘continuation probability’ () denotes how likely a user is to continue to browse to the next statement after having seen one. This can alternatively be framed in terms of the ‘weight’ (), which refers to the probability of a user reaching each step of the sequence. The ‘last probability’ (), indicates whether a given statement is the last one to be read before aborting the search.
Progression indicates that the assumptions made by the framework are applicable in the first place, requiring a sequential process. Decay is encoded by a diminishing attention, relating to continuation probability and weight (/), while saturation relates to the last probability (). In sum, this allows to operationalize the reading model as a monotonically decreasing weight function over statements, discounting the contribution of statements occurring later in the response. This induces a corresponding response-level document organization where the most important pieces of information come first and are then followed by increasingly insignificant details, like the inverted pyramid scheme of news articles (Pöttker, 2003).
3.4.3. Accumulation Model for Generative IR

### v3

@@ 3.1. Preliminary Considerations
Traditional search results (list SERPs) are ranked lists of documents, each typically referenced by a linked title, snippet, and URL. In generative IR, instead, the search result is a textual response (text SERP), i.e., a sequence of statements, each optionally referenced to sources of evidence. A statement can be any consecutive passage of text, ranging from phrases to sentences or even longer paragraphs. In this context, we consider statements as atomic in the sense that we disregard the nesting of statements of different lengths, and in the sense that statements support claims that are pertinent to the user’s information need—comparable to the concept of ‘atomic/semantic content units’ (Liu et al., 2023a; Nenkova et al., 2007) in summarization evaluation, or ‘information nuggets’ / ‘retrieval units’ in traditional IR (Dang and Lin, 2007; Sakai et al., 2011; Sakai, 2023; Chen et al., 2023). A statement can be referenced to none, one, or more sources in form of explicit links to web documents containing the information on which the generated statement is based and by which it is grounded. In this paper, we consider the evaluation to be ad hoc, i.e., based on a single query without search session-based or conversational elements.
@@
To estimate the effectiveness of list SERP-based retrieval systems, offline evaluation following the Cranfield paradigm (Cleverdon, 1967) is a de facto standard in IR research. The users’ satisfaction with the results for a given topic (query) is estimated by deriving effectiveness scores based on judging a pool of documents returned by the evaluated systems (Sanderson, 2010). The pools often are also reused later to evaluate new search systems by checking whether their retrieved results previously were judged—and simplistically assuming non-relevance for the results without previous judgments (Fröbe et al., 2023). However, as the output “documents” of generative retrieval systems may be novel every time, simply assuming non-relevance would not lead to helpful evaluation results. Instead, more sophisticated transfer methods are required to adapt offline evaluation to generative retrieval. Besides offline evaluation, generative retrieval systems could also be evaluated in an online fashion (Sallam et al., 2023). Online evaluation does not rely on previous judgments but tries to estimate the output of some system by collecting explicit or implicit user feedback (Kelly, 2009) like user satisfaction ratings or clicks. This form of evaluation increases the manual effort, often happens in uncontrolled setups, may be expensive and time-consuming to conduct, and is challenging to replicate, repeat, and reproduce (Renaud and Azzopardi, 2012). To mitigate these issues especially in an academic setting with limited access to human user data, some studies suggested user simulation to analyze (interactive) information systems (Maxwell et al., 2015; Maxwell and Azzopardi, 2016; Câmara et al., 2022; Zerhoudi et al., 2022). However, simulated users cannot yet compete with “real” human feedback. Recently, fully automatic evaluations, where the output of one system is judged by another, has been proposed as a possible way forward (Liu et al., 2023b; Yue et al., 2023). But judging the output of generative models by means of other models has itself already been criticized (Sakai, 2023; Bauer et al., 2023; Faggioli et al., 2023).
3.2. Steps of the Information Search Process
@@
To derive suitable evaluation objectives for generative ad hoc retrieval, we consider the general user side search process for which Vakkari (Vakkari, 2016) has suggested to differentiate four steps: search formulation, source selection, source interaction, and information synthesis (Figure 4, top row). Interestingly, each of these steps can be mapped to capabilities of generative retrieval systems.
@@
First, during formulation, the user comes up with a specific query that expresses their information need. This is no different in generative retrieval systems, though what is called a ‘query’ in IR is often called a ‘prompt’ for generative systems. To avoid confusion, we stick to the term ‘query’. Still, in this paper, we leave the formulation step entirely to the user who may iteratively adapt their search formulation. Yet, we do acknowledge that formulation may also be framed as a system task with the goal of enhancing the users’ original query with more context or prompt templates, akin to query suggestion and query expansion in traditional retrieval. Second, during selection, traditionally, the user is presented with a result list possibly containing surrogates like snippets that help to assess whether some result aligns with the user’s information need and should be selected for further inspection. In generative retrieval, the selection step corresponds to the system selecting sources that contain potentially relevant information. Third, during interaction, traditionally, the user analyzes the content of the selected results more deeply to extract and structure the relevant information that addresses the knowledge gap underlying the user’s information need. In generative retrieval, this step also rather is on the system side by, for instance, attending the generation to previously retrieved pieces of information. Finally, during synthesis, traditionally, the user assembles the search outcome by combining relevant information from their interacted sources. In generative retrieval, synthesis corresponds to the model’s inference and generation of the response text from the selected sources. Just like for human users, interaction and synthesis may commence concurrently.
@@ 3.3. Evaluation Objectives
For each step of the search process, we define a corresponding generative retrieval-oriented evaluation objective. The objectives are not meant as evaluation steps, but rather as potential targets when evaluating a generative retrieval system as a whole.
@@
Corresponding to formulation are evaluation aspects related to a model’s input prompt like preciseness (Does the prompt target the specific desired outcome?), ambiguity (Is the prompt unambiguous, targeting only the desired outcome?), or contextuality (Does the prompt provide sufficient context to delineate the information need?). While formulation is an important step to evaluate, it is out of the scope of our paper, as the formulation step in our setting is left to the user and as there already is extense work on prompt engineering (Shin et al., 2020; Reynolds and McDonell, 2021; Gao et al., 2021; Liu and Chilton, 2022; Sorensen et al., 2022; White et al., 2023; Yang et al., 2023).
@@
Corresponding to selection are evaluation aspects related to the retrieved sources from which a generative system draws its information. These sources (but also any relevant information that was not retrieved) directly impact the quality of the generated response. Therefore, the retrieval objective covers the assessment of a system’s ability to identify relevant (aligning with the users’ information need), diverse (covering a variety of information), informative (containing valuable information), and correct (providing accurate information) sources from a collection.
@@
Corresponding to interaction are evaluation aspects related to a generative retrieval model’s ability to attend to source documents as evidence in response generation. Yet, such grounded text generation may suffer from confabulations / hallucinations of broadly two types (Maynez et al., 2020): intrinsic confabulations (the model wrongly modifies information from the sources) and extrinsic confabulations (the model generates information not present in the sources). As both types can negatively impact the quality of a generated response (Maynez et al., 2020; Lux et al., 2020), the grounding objective covers the assessment of a system’s ability to correlate its generated output with information from source documents. This includes the ability to identify relevant information in the sources, to paraphrase information (restate some information correctly), and to establish consistency (not produce contradictions to other sources).
@@
Corresponding to synthesis are evaluation aspects related to a model’s ability to condense relevant information from multiple sources into a single answer. Resembling multi-document summarization, the presentation objective covers the assessment of an answer’s conciseness (at a level of granularity sensible given the topic and user (Dang, 2005)), coherence (uniform writing style in the answer), and accessibility (written in an understandable way; again, dependent on the user).
@@ 3.4. Components of the User Model
Developing a user model for generative IR is challenging. The traditional user models were focusing on list SERPs and might thus not apply to text SERPs, as, for instance, the search process steps of selection and interaction are undertaken by the system instead. Additionally, little to no user behavior data on text SERPs are available in the academic context (e.g., A/B tests or laboratory studies) to base model validation or development on. To contribute a user model for generative IR, we thus extrapolate from established evaluation practices in IR and related fields like question answering or summarization. We follow the considerations of Carterette (2011), who argues that a user model in IR should include three distinct (sub-)models: (1) a utility model that induces a gain function by capturing how each result provides utility to a user, (2) a browsing model that induces a discount function by capturing how a user interacts with the results, and (3) an accumulation model that combines the individual gain and discount values by capturing how individual utility is aggregated.
3.4.1. A Utility Model for Generative IR
Surveying the literature on evaluation in IR and in related fields, we identified ten utility dimensions applicable to generative ad hoc retrieval. Figure 5 shows the dimensions grouped into five categories (coherence, coverage, consistency, correctness, and clarity) with color-coded corresponding evaluation objectives and indicated granularity from which gain is obtained (statement level: from an individual statement in the response; response level: from the response as a whole).
@@ 3.4.1. A Utility Model for Generative IR
Coherence is a response-level dimension of utility referring to the presentation objective and involving the aspects of statement arrangement that should form a narrative without contradictions (Radev and McKeown, 1998; Shah et al., 2021) (i.e., logical coherence: Is the response well-structured?) and of the writing style that should yield readable and engaging responses (Jin et al., 2020; Capra and Arguello, 2023) (i.e., stylistic coherence: Does the response have a uniform style of speech?).
@@
Coverage is a response-level dimension of utility referring to the retrieval objective and measuring how well a user’s information need is treated by the returned information; it can be subdivided into (Cambazoglu et al., 2021) broad coverage (i.e., whether the response covers diverse information (Zheng et al., 2012)), and deep coverage (i.e., whether the response provides in-depth and highly informative content (Maxwell et al., 2017)).
@@
Commonly observed problems with source-based text generation are inconsistencies between the sources and parts of the generated text (Huang et al., 2021) but also inconsistencies between the statements within a response. We refer to the first problem as external consistency, which is a statement-level dimension of utility involving the assessment of the consistency between a statement and its source document(s) to ensure that the generated text aligns in terms of content and context (Maynez et al., 2020; Yue et al., 2023; Sakai, 2023) (i.e., Is the statement accurately conveying from the sources?). External inconsistencies are often introduced through model confabulations / hallucinations (Ji et al., 2023) but they should be distinguished from factual correctness, as external consistency only assesses the alignment of a statement with the sources, and not with some objective truth. To the second consistency problem, we refer to as internal consistency, which is a response-level dimension of utility involving the assessment of the consistency between the responses’ individual statements to ensure no contradictions (Nishino et al., 2019; Sakai, 2023; Capra and Arguello, 2023). It should be noted that this does not mean that different conflicting perspectives on a topic can not be reflected in the response, however, these should then be explained. Both notions of consistency refer to the grounding objective.
@@
Correctness is a statement-level dimension of utility referring to the retrieval objective and measuring to which degree the information provided in the response is factual, reliable, and addressing the user’s information needs. We subdivide correctness into factual and topical correctness. The former captures the degree to which a statement reproduces information that can be assumed as objectively true. Yet, outside of small-scale domain-specific evaluation studies (Sallam et al., 2023) fact-checking remains a challenge (Nakov et al., 2021) and is thus often reduced to a simpler approach, framing it in terms of verifiability (Liu et al., 2023b), not truth. Here, the main requirement is that a piece of information can be attributed to a reliable reference (Wikimedia Foundation, 2023; Yue et al., 2023) (i.e., Does the statement state things that are verifiable?). Topical correctness captures whether a statement aligns with the user’s information need (Maddalena et al., 2017; Yang, 2017; Roitero et al., 2018) (i.e., Does the statement state things within the scope of the user’s information need?).
@@
The response of a generative retrieval system should be expressed in a clear and understandable way (Zhu et al., 2009; Sameki et al., 2016). This, on the one hand, comprises language clarity: concise (Dang, 2005; Sakai, 2023), comprehensible (Cambazoglu et al., 2021), lexically and grammatically correct, and user-accessible responses. Note that language clarity does not reflect fluency, which is assumed already at human-level for model-generated text (Sakai, 2023), but rather the response being in the appropriate language register. For example, a technical query might warrant an academic style of writing in the response, while a joke question might afford a more jovial tone. On the other hand, clarity also comprises content clarity: in order to make a response explainable, the way a statement is written should always clearly communicate the most salient information (Schuff et al., 2022) and where it stems from (Nourani et al., 2019). Both notions of clarity refer to the presentation objective at the statement level.
3.4.2. A “Browsing” Model for Generative IR
@@
For list SERPs, user interaction is modeled by a set-based or a ranking-based browsing model. In set-based browsing users are assumed to indiscriminately examine all retrieved documents (e.g., for systematic reviews), while in ranking-based browsing users are assumed to traverse the retrieved documents by rank, stopping when either their information need is fulfilled or the search is aborted (Carterette, 2011) (e.g., web search). Aborting a search is usually motivated by the information need being satisfied or the effort being too high to justify continuing to browse. Yet, in generative IR, the selection and interaction steps of the search process are undertaken by the system, so that the user only has to read the (often short) generated text. This reduces the effect of effort-based stopping criteria, with most users only aborting their search when their knowledge gap is fulfilled or when the response is deemed insufficient. This is neither really set-based, as reading the response from the beginning and early stopping might occur, nor traditionally ranking-based, as aborting the search is not motivated by effort but rather by search (dis-)satisfaction.
@@
Instead of a browsing model, we thus propose a reading model for generative IR to reflect the attention a user places on the response statements while reading. But as there are no dedicated studies on reading behavior for generative IR yet, we turn to related work on reading behavior for document comprehension. In a literature survey, we identified six characteristics from which we deem three as appropriate for our reading model: progression (Buscher et al., 2012; Li et al., 2018, 2019; Zheng et al., 2019; Wu et al., 2023) implies that users parse a document sequentially (i.e., reading the statements in their textual order), decay (Frey et al., 2013; Li et al., 2018, 2019; Zheng et al., 2019; Wu et al., 2023) implies that the reading attention diminishes over the span of a text, and saturation (Li et al., 2018, 2019) implies that users abort when they have read enough to satisfy their information need.
@@
Besides these three characteristics, we deem three others as superfluous for our proposed reading model. First, perceived relevance may be heightened following a relevant statement (Li et al., 2018; Zheng et al., 2019) but we adopt the same restriction as static browsing models for ad hoc retrieval evaluation and neglect inter-statement effects (Moffat et al., 2013, 2015). Second, although reading attention may be highest around query terms (Li et al., 2018; Zheng et al., 2019), our statement- and response-level utility granularities render per-token effects rather constant. Third, although users may skip non-relevant content during reading (Li et al., 2018; Gwizdka, 2014; Buscher et al., 2012), we ignore this effect in the reading model as non-relevant statements will receive zero utility anyway.
Altogether, our proposed reading model thus reflects a sequential reading with decaying attention and early stopping when saturated. These properties can easily be related to the framework (Moffat et al., 2017) of browsing models for list SERPs. Sequential reading indicates that the framework’s assumption of a sequential process applies, decay (diminishing attention) is related to the framework’s conditional continuation probability and weight (probability of a user reaching some step of a sequence), and early stopping (saturation) is related to the framework’s probability that indicates whether an item is the last one before aborting a search. Our proposed reading model can thus be operationalized as a monotonically decreasing weight function over statements that discounts the contribution of later statements in a response. At the same time, this directly induces a response structuring approach of putting the most important pieces of information first followed by less important details—similar to the inverted pyramid scheme of news articles (Pöttker, 2003).
3.4.3. An Accumulation Model for Generative IR